OpenStudy (anonymous):

What is the vertex of the graph of y = 2x^2 − 3x − 1?

OpenStudy (anonymous):

y=2(x^2-3/2x)-1 y=2(x-3/4)^2-9/8-1 y+17/8=2(x-3/4)^2. Hence vertex is at 3/4,-17/8